# The Formality of the Goldman–Turaev Lie Bialgebra on a Closed Surface

### v1

@@ Abstract
We reformulate the Kashiwara–Vergne groups and associators in higher genera, introduced in Alekseev–Kawazumi–Kuno–Naef, in terms of non-commutative connections using the tools developed in a previous paper. As the main result, the case of closed surfaces is dealt with to determine the pro-nilpotent automorphism group of the associated graded of the Goldman–Turaev Lie bialgebra.
@@ 1. Introduction
The defining relations of the KV associators and the KV groups have a higher genus analogue. Alekseev–Kawazumi–Kuno–Naef [AKKN23] shows that, in the genus 00 0 case, the formality problem of the Goldman–Turaev Lie bialgebra is essentially equivalent to the KV problem; this perspective immediately allows us to define the higher genus KV associators as the solutions to the formality problem and the KV groups as the (pro-unipotent) automorphism group of the (graded) Goldman–Turaev Lie bialgebra. On the other hand, we have higher genus analogues of the Drinfeld associators. For genus 11 11, they are called elliptic associators introduced by Enriquez [Enr14], and their relation to the corresponding KV problem is discussed in [AKKN23]. For an arbitrary genus, several generalisations of the GT groups and the Drinfeld associators are proposed: [Gon20] by Gonzalez, [Fel21] by Felder, and Section 5 of [CIW19] by Campos–Idrissi–Willwacher. They all agree in genus 11 11, but the relation between them and the KV associators for genus $\geq 2$ is still an open question.
@@
Getting back to the KV problem in higher genera, the existence problem itself is completely solved in [AKKN23], and the key step in the proof is the factorisation of the Turaev cobracket in two parts: a Hamiltonian flow and a non-commutative divergence. In this paper, we solve the formality problem on a closed surface using the factorisation of the Turaev cobracket obtained in the author’s previous paper [Tan24b]. We start with a reformulation of the KV groups and the space of solution $\mathrm{SolKV}$ for a connected oriented compact surface $\Sigma$ with non-empty boundary in terms of connections in non-commutative geometry and later see the case of closed surfaces based on the observation.
@@
Let $\mathbb{K}$ be a field of characteristic zero, $\pi=\pi_{1}(\Sigma)$ the fundamental group of $\Sigma$ and $H=H_{1}(\Sigma;\mathbb{K})$ the first homology group. In their formulation, the KV groups are described in terms of group 11 11-cocycles on the group $\operatorname{tAut}_{\mathrm{Hopf}}(\widehat{\mathbb{K}\pi})$ of tangential automorphisms on the complete Hopf algebra $\widehat{\mathbb{K}\pi}$, and the graded counterpart $\operatorname{tAut}_{\mathrm{Hopf}}(\hat{T}(H))$ on the completed tensor algebra $\hat{T}(H)$. We first see that a tangential automorphism of $\widehat{\mathbb{K}\pi}$ is interpreted as a derivation of the complete Hopf groupoid (in the sense of Fresse [Fre17]) $\widehat{\mathbb{K}\mathscr{G}}$ associated with the fundamental groupoid $\mathscr{G}$ of $\Sigma$, and similarly for the graded counterpart $\hat{\mathscr{T}}$. We also have their subcategories $\partial\widehat{\mathbb{K}\mathscr{G}}$ and $\partial\hat{\mathscr{T}}$ generated by boundary loops.
@@
Next, we define connections $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ and $\nabla^{\prime}\!\!_{H,\mathsf{fr}}$ on some $\mathbb{K}\mathscr{G}$-module and $\hat{\mathscr{T}}$-module, respectively, depending on a framing $\mathsf{fr}$ on $\Sigma$ and a free-generating system $\mathcal{C}$ of $\pi$. Recall that a connection on a $A$-module $N$ is a linear map $\nabla\colon N\to\Omega^{1}A\otimes_{A}N$ satisfying the Leibniz rule, where $\Omega^{1}A$ is the space of non-commutative 11 11-forms. Then, the value of those 11 11-cocycles on an automorphism $G$ is written as (the trace of) the difference of connections by the action of $G$: $\operatorname{Tr}(\operatorname{Ad}_{G}\!\nabla-\nabla)$ where $\nabla$ is either of the two above. This takes its value in the space of de Rham 1-forms $\operatorname{DR}^{1}\!A$, a quotient space of $\Omega^{1}A$. We have the exterior derivative $d$, which sends an element of the cyclic quotient $|A|=A/[A,A]$ to $\operatorname{DR}^{1}\!A$. One of the main results is the following.
@@
Here, the subscript + indicates the pro-unipotent part of each space. In particular, the Lie algebra corresponding to $\mathrm{KRV}^{\mathsf{fr}}_{(g,n+1)}$ is given by
@@
Here, $\hat{L}(H)_{\omega}$ is some quotient of the completed free Lie algebra over $H$, $\nabla^{\prime}_{\bullet,H}$ is a connection on a $\hat{L}(H)_{\omega}$-module and $\bar{\Delta}_{\omega}\colon\hat{T}(H)_{\omega}\to\hat{T}(H)_{\omega}\hat{\otimes}\hat{T}(H)_{\omega}$ is the reduced coproduct: $\bar{\Delta}_{\omega}(x)=\Delta_{\omega}(x)-x\otimes 1-1\otimes x$.
The space $\operatorname{Ker}(|\bar{\Delta}_{\omega}|)$ deserves some description, so we computed its lower degree parts at the end of the paper.
Organisation of the paper. In Sections 2 and 3, we recall the definition of Goldman–Turaev Lie bialgebra, the formality problem, and main results in [AKKN23]. Section 4 is occupied with the reformulation of the KV groups. In Section 5, we construct a basis of $|T(H)_{\omega}|$ for later sections. The formality problem on a closed surface is dealt with in Section 6. Section 7 is devoted to the computation of the kernel of the reduced coproducts.
Acknowledgements. The author thanks Yusuke Kuno for a discussion on Section 7 and Nariya Kawazumi for thoroughly reading a draft.
@@ 2. The Goldman–Turaev Lie Bialgebra
In this section, we recall the Goldman–Turaev Lie bialgebra and its associated graded following [AKKN23]. Let $g\geq 0$, $n\geq-1$ and $\Sigma=\Sigma_{g,n+1}$ a connected compact oriented surface of genus $g$ and $n+1$ boundary components $\partial_{0}\Sigma,\dotsc,\partial_{n}\Sigma$. Consider the fundamental group $\pi=\pi_{1}(\Sigma,*)$ of $\Sigma$ and the group algebra $\mathbb{K}\pi$. We require the base point * to be taken on $\partial_{0}\Sigma$ if $n\geq 0$. For a $\mathbb{K}$-algebra $A$, we denote the cyclic quotient $A/[A,A]$ by $|A|$. Then, $|\mathbb{K}\pi|=\mathbb{K}\pi/[\mathbb{K}\pi,\mathbb{K}\pi]$ is identified with the $\mathbb{K}$-vector space spanned by the homotopy set of free loops on $\Sigma$.
@@
where $\mathrm{sign}(p;\alpha,\beta)$ is the local intersection number of the velocity vectors of $\alpha$ and $\beta$ at $p$ with respect to the orientation of the surface, and $\alpha*_{p}\beta$ is the free loop obtained by first traversing $\alpha$ starting at $p$ then $\beta$.
@@
Assume $n\geq 0$. We define the Kawazumi–Kuno action $\sigma\colon|\mathbb{K}\pi|\to\operatorname{End}_{\mathbb{K}}(\mathbb{K}\pi)$ by, for generically immersed free loops $\alpha$ and $\beta$ on $\Sigma$,
@@
  • Consider a morphism of $\mathbb{K}$-algebras
@@
  • A filtration on $\operatorname{Der}_{\mathbb{K}}(\mathbb{K}\pi)$ is the induced one: denoting the filtration degree $\geq k$ part of $\mathbb{K}\pi$ by $F^{k}\mathbb{K}\pi$, $u\colon\mathbb{K}\pi\to\mathbb{K}\pi$ has filtration degree $\geq p$ if and only if $u(F^{k}\mathbb{K}\pi)\subset F^{k+p}\mathbb{K}\pi$ for all $k$.
@@
For $n=-1$, a filtration on $|\mathbb{K}\pi_{1}(\Sigma_{g,0})|$ is defined by the push-out along $|\mathbb{K}\pi_{1}(\Sigma_{g,1})|\twoheadrightarrow|\mathbb{K}\pi_{1}(\Sigma_{g,0})|$ by capping the unique boundary component by a disk. The rest is analogously defined.
@@ 3. The Formality and the Kashiwara–Vergne Problem
The formality problem for the Goldman–Turaev Lie algebra asks if it is isomorphic to its associated graded as a Lie bialgebra and, if so, to determine the set of such isomorphisms. In the case of a surface with non-empty boundary, it is completely solved in [AKKN23]: the set SolKV of such isomorphisms is characterised by two equations (KVI) and (KVII). In the special case of $\Sigma=\Sigma_{0,3}$, the problem is, surprisingly, equivalent to the Kashiwara–Vergne problem in Lie theory, and the automorphism group of the Lie bialgebra and that of the associated graded are given by the KV and KRV groups. This section summarises definitions of these spaces and results in [AKKN23].
@@
  • $\hat{T}(H)$ is regarded as a $U\!\hat{L}(H)$-module by the left multiplication. For $w=x_{i},y_{i},z_{j}$, $d_{w}\colon\hat{L}(H)\to\hat{T}(H)$ is a continuous Lie algebra 11 11-cocycle specified by $d_{w}(w^{\prime})=\delta_{ww^{\prime}}$ for $w^{\prime}=x_{i},y_{i},z_{j}$ using Kronecker’s delta.
@@
    $\operatorname{div}_{x,y,z}\colon\operatorname{tDer}(\hat{L}(H))\to\operatorname{Der}(\hat{L}(H))\xrightarrow{\operatorname{div}_{x,y,z}}|\hat{T}(H)|.$
@@ 4.1 Tangential Derivations
Proof. Suppose we are given a tangential derivation $(u,u_{1},\dotsc,u_{n})$. We construct a $\mathbb{K}$-linear category derivation $f$. For a loop $\alpha$ based at $*_{0}$, we set $f(\alpha)=u(\alpha)$. For a path $\delta$ from $*_{j}$ to $*_{0}$ such that $\delta\gamma_{j}\delta^{-1}$ is a $j$-th boundary loop, we set $\delta^{-1}f(\delta)=u_{j}$. This is well-defined: for another such path $\delta^{\prime}$, the loop $\delta^{\prime-1}\delta$ is based at $*_{0}$ and we have $(\delta^{\prime-1}\delta)\gamma_{j}=\gamma_{j}(\delta^{\prime-1}\delta)$. Since $\mathscr{G}$ is a free groupoid, we have $\delta^{\prime-1}\delta=\gamma_{j}^{-m}$ for some $m\in\mathbb{Z}$. We compute
@@ 4.2 Connections on Modules
A $\mathbb{K}$-Hopf groupoid structure of the groupoid algebra $\mathbb{K}\mathscr{G}$ is given by the following: a counital coassociative (cocommutative) coalgebra structure on the Hom-spaces given by, for $v,w\in\mathrm{Ob}(\mathscr{G})=V$ and $x\in\mathscr{G}(v,w)$,
@@
The completion $\widehat{\mathbb{K}\mathscr{G}}$ and its associated graded $\hat{\mathscr{T}}$ are automatically (complete) Hopf groupoids. We set
@@
The proof is analogous to Proposition 3.8 in [Tan24b] and Lemma 5.6 in [Tan24a].∎
@@
Take a free-generating system of $\mathscr{G}$ as in Figure 2 of [Tan24a]. We define a connection $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}\colon\mathscr{N}\to\Omega^{1}\mathbb{K}\mathscr{G}\otimes_{\mathbb{K}\mathscr{G}}\mathscr{N}$ by
@@
Extend $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ to the completion $\hat{\mathscr{N}}$ by continuity, which is a $\widehat{\mathbb{K}\mathscr{G}}$-module. A derivation action of $f\in\operatorname{Der}_{\mathrm{Hopf}}(\widehat{\mathbb{K}\mathscr{G}})$ on the module $\hat{\mathscr{N}}$ is given by a restriction $f|_{\operatorname{Ker}(s)}$.
@@
Let $\hat{\mathscr{N}}_{\mathrm{gr}}=\operatorname{Ker}(s\colon\hat{\mathscr{T}}\to\mathbb{K}V)$, which is also dualisable over $\hat{\mathscr{T}}$. We define a connection $\nabla^{\prime}\!\!_{H,\mathsf{fr}}$ on $\hat{\mathscr{N}}_{\mathrm{gr}}$ by
@@
This is an extension of the flat connection $\nabla^{\prime}\!\!_{z}$ defined in Section 6 of [Tan24b]. Also this is the associated graded of $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ above.
@@
Before the proof, we look at a counterpart of the connection $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ on a $\mathbb{K}\mathscr{G}$-bimodule $\Omega^{1}\mathbb{K}\mathscr{G}$. For a $\mathbb{K}$-algebra $A$, we denote the copy of an element $a\in A$ in $A^{\mathrm{op}}$ by $\bar{a}$. We denote by $A^{\mathrm{e}}$ the enveloping algebra of $A$.
@@
We define a connection $\nabla\!_{\mathcal{C},\mathsf{fr}}\colon\Omega^{1}\mathbb{K}\mathscr{G}\to\Omega^{1}\mathbb{K}\mathscr{G}^{\mathrm{e}}\otimes_{\mathbb{K}\mathscr{G}^{\mathrm{e}}}\Omega^{1}\mathbb{K}\mathscr{G}$ on a $\mathbb{K}\mathscr{G}$-bimodule $\Omega^{1}\mathbb{K}\mathscr{G}$ by
@@
First, the source map corresponds to a left $A$-module map $s\colon A\to\mathbb{K}V$, where the corresponding $A$-module to $\mathbb{K}V$ is denoted again by $\mathbb{K}V$. In addition, $\mathscr{N}$ corresponds to $N=\operatorname{Ker}(s\colon A\to\mathbb{K}V)$. Next, consider the functor

### v2

@@ Abstract
We reformulate the Kashiwara–Vergne groups and associators in higher genera, introduced in Alekseev–Kawazumi–Kuno–Naef, in terms of non-commutative connections using the tools developed in a previous paper. As the main result, the case of closed surfaces is dealt with to determine the pro-unipotent automorphism group of the associated graded of the Goldman–Turaev Lie bialgebra.
@@ 1. Introduction
The defining relations of the KV associators and the KV groups have a higher genus analogue. Let $\mathbb{K}$ be a field of characteristic zero and $\pi=\pi_{1}(\Sigma)$ the fundamental group of a compact oriented connected surface $\Sigma$. Then, the trace space $|\mathbb{K}\pi|=\mathbb{K}\pi/[\mathbb{K}\pi,\mathbb{K}\pi]$ is endowed with the structure of the Goldman–Turaev Lie bialgebra. Alekseev–Kawazumi–Kuno–Naef [AKKN23] shows that, in the genus 00 0 case, the formality problem of the Goldman–Turaev Lie bialgebra, which asks if the completion of this Lie bialgebra under a suitable filtration is isomorphic to its associated graded (see Section 3 for the detail), is essentially equivalent to the KV problem; this perspective immediately allows us to define the higher genus KV associators as the solutions to the formality problem and the KV groups as the (pro-unipotent) automorphism group of the (graded) Goldman–Turaev Lie bialgebra. We remark that, also in the genus 00 0 case, a direct construction of a formality isomorphism from a Drinfeld associator by means of the LMO functor is originally given in [Mas18] by Massuyeau. On the other hand, we have higher genus analogues of the Drinfeld associators. For genus 11 11, they are called elliptic associators introduced by Enriquez [Enr14], and their relation to the corresponding KV problem is discussed in [AKKN23]. For an arbitrary genus, several generalisations of the GT groups and the Drinfeld associators are proposed: [Gon20] by Gonzalez, [Fel21] by Felder, and Section 5 of [CIW19] by Campos–Idrissi–Willwacher. They all agree in genus 11 11, but the relation between them and the KV associators for genus $\geq 2$ is still an open question.
@@
Getting back to the KV problem in higher genera, the existence problem itself is completely solved in [AKKN23], and the key step in the proof is the factorisation of the Turaev cobracket in two parts: a Hamiltonian flow and a non-commutative divergence. In this paper, we solve the (non-)uniqueness part of the formality problem on a closed surface: we determine the pro-unipotent part of the automorphism group of the Goldman–Turaev Lie bialgebra on a closed surface using the factorisation of the Turaev cobracket obtained in the author’s previous paper [Tan24b]. We start with a reformulation of the KV groups and the space of solution $\mathrm{SolKV}$ for a surface $\Sigma$ with non-empty boundary in terms of connections in non-commutative geometry and later see the case of closed surfaces based on the observation.
@@
Let $H=H_{1}(\Sigma;\mathbb{K})$ be the first homology group of the surface $\Sigma$. In their formulation, the KV groups are described in terms of group 11 11-cocycles on the group $\operatorname{tAut}_{\mathrm{Hopf}}(\widehat{\mathbb{K}\pi})$ of tangential automorphisms on the complete Hopf algebra $\widehat{\mathbb{K}\pi}$, and the graded counterpart $\operatorname{tAut}_{\mathrm{Hopf}}(\hat{T}(H))$ on the completed tensor algebra $\hat{T}(H)$. We first see that a tangential automorphism of $\widehat{\mathbb{K}\pi}$ is interpreted as a derivation of the complete Hopf groupoid (in the sense of Fresse [Fre17]) $\widehat{\mathbb{K}\mathscr{G}}$ associated with the fundamental groupoid $\mathscr{G}$ of $\Sigma$, and similarly for the graded counterpart $\hat{\mathscr{T}}$. We also have their subcategories $\partial\widehat{\mathbb{K}\mathscr{G}}$ and $\partial\hat{\mathscr{T}}$ generated by boundary loops.
@@
Next, we define connections $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ and $\nabla^{\prime}\!\!_{H,\mathsf{fr}}$ on some $\mathbb{K}\mathscr{G}$-module and $\hat{\mathscr{T}}$-module, respectively, depending on a framing $\mathsf{fr}$ on $\Sigma$ and a free-generating system $\mathcal{C}$ of $\pi$. Recall that a connection on an $A$-module $N$, where $A$ is a $\mathbb{K}$-algebra, is a $\mathbb{K}$-linear map $\nabla\colon N\to\Omega^{1}A\otimes_{A}N$ satisfying the Leibniz rule, where $\Omega^{1}A$ is the space of non-commutative 11 11-forms. Then, the value of those 11 11-cocycles on an automorphism $G$ is written as (the trace of) the difference of connections by the action of $G$: $\operatorname{Tr}(\operatorname{Ad}_{G}\!\nabla-\nabla)$ where $\nabla$ is either of the two above. This takes its value in the space of de Rham 1-forms $\operatorname{DR}^{1}\!A$, a quotient space of the space $\Omega^{1}A$ of non-commutative 11 11-forms. We have the exterior derivative $d$, which sends an element of the trace space $|A|=A/[A,A]$ to $\operatorname{DR}^{1}\!A$. One of the main results is the following.
@@
Here, the superscript + indicates the pro-unipotent part of each space. In particular, the Lie algebra corresponding to $\mathrm{KRV}^{\mathsf{fr}}_{(g,n+1)}$ is given by
@@
Here, $\hat{L}(H)_{\omega}$ is the quotient of the completed free Lie algebra over $H$ by the element $\omega\in H^{\otimes 2}$ representing the symplectic structure of $H$, $\nabla^{\prime}_{\bullet,H}$ is a connection on a $\hat{L}(H)_{\omega}$-module and $\bar{\Delta}_{\omega}\colon\hat{T}(H)_{\omega}\to\hat{T}(H)_{\omega}\hat{\otimes}\hat{T}(H)_{\omega}$ is the reduced coproduct: $\bar{\Delta}_{\omega}(x)=\Delta_{\omega}(x)-x\otimes 1-1\otimes x$.
The space $\operatorname{Ker}(|\bar{\Delta}_{\omega}|)$ deserves some description, so we compute it in small degrees at the end of the paper.
Organisation of the paper. In Sections 2 and 3, we recall the definition of the Goldman–Turaev Lie bialgebra, the formality problem, and main results in [AKKN23]. Section 4 is occupied with the reformulation of the KV groups. In Section 5, we construct a basis of $|T(H)_{\omega}|$ for later sections. The formality problem on a closed surface is dealt with in Section 6. Section 7 is devoted to the computation of the kernel of the reduced coproducts.
Acknowledgements. The author thanks Yusuke Kuno for a discussion on Section 7, Nariya Kawazumi for thoroughly reading the draft and Geoffrey Powell for many comments and improvements on the draft.
@@ 2. The Goldman–Turaev Lie Bialgebra
In this section, we recall the Goldman–Turaev Lie bialgebra and its associated graded following [AKKN23]. Let $g\geq 0$, $n\geq-1$ and $\Sigma=\Sigma_{g,n+1}$ a connected compact oriented surface of genus $g$ and $n+1$ boundary components $\partial_{0}\Sigma,\dotsc,\partial_{n}\Sigma$. Consider the fundamental group $\pi=\pi_{1}(\Sigma,*)$ of $\Sigma$ and the group algebra $\mathbb{K}\pi$. We require the base point * to be taken on $\partial_{0}\Sigma$ if $n\geq 0$. For a $\mathbb{K}$-algebra $A$, we denote the trace space $A/[A,A]$ by $|A|$. Then, $|\mathbb{K}\pi|=\mathbb{K}\pi/[\mathbb{K}\pi,\mathbb{K}\pi]$ is identified with the $\mathbb{K}$-vector space spanned by the homotopy set of free loops on $\Sigma$.
@@
where $\mathrm{sign}(p;\alpha,\beta)$ is the local intersection number of the velocity vectors $\dot{\alpha}_{p}$ and $\dot{\beta}_{p}$ at $p$ with respect to the orientation of the surface, and $\alpha*_{p}\beta$ is the free loop obtained by first traversing $\alpha$ starting at $p$ then $\beta$.
@@
Assume $n\geq 0$. Let $\alpha$ be a free loop and $\beta$ a based loop on $\Sigma$, both generically immersed. We define the Kawazumi–Kuno action $\sigma\colon|\mathbb{K}\pi|\to\operatorname{End}_{\mathbb{K}}(\mathbb{K}\pi)$ by
@@
  • Consider the morphism of $\mathbb{K}$-algebras
@@
  • The filtration on $\operatorname{Der}_{\mathbb{K}}(\mathbb{K}\pi)$ is the induced one: denoting the filtration degree $\geq k$ part of $\mathbb{K}\pi$ by $F^{k}\mathbb{K}\pi$, $u\colon\mathbb{K}\pi\to\mathbb{K}\pi$ has filtration degree $\geq p$ if and only if $u(F^{k}\mathbb{K}\pi)\subset F^{k+p}\mathbb{K}\pi$ for all $k$.
@@
For $n=-1$, the filtration on $|\mathbb{K}\pi_{1}(\Sigma_{g,0})|$ is defined by the push-out along $|\mathbb{K}\pi_{1}(\Sigma_{g,1})|\twoheadrightarrow|\mathbb{K}\pi_{1}(\Sigma_{g,0})|$ by capping the unique boundary component by a disk. The rest is analogously defined.
@@ 3. The Formality and the Kashiwara–Vergne Problem
The formality problem for the Goldman–Turaev Lie algebra asks if it is isomorphic to its associated graded as a Lie bialgebra and, if so, to determine the set of such isomorphisms. More precisely, a solution to the formality problem is a completed Hopf algebra isomorphism
$\theta\colon\widehat{\mathbb{K}\pi}\to\mathrm{gr}(\widehat{\mathbb{K}\pi})$
such that $\mathrm{gr}(\theta)=\operatorname{id}$ and induces an isomorphism of Lie bialgebras $|\widehat{\mathbb{K}\pi}|\to|\mathrm{gr}(\widehat{\mathbb{K}\pi})|$.
In the case of a surface with non-empty boundary, it is completely solved in [AKKN23]: the set SolKV of such isomorphisms is characterised by two equations (KVI) and (KVII). In the special case of $\Sigma=\Sigma_{0,3}$, the problem is, surprisingly, equivalent to the Kashiwara–Vergne problem in Lie theory, and the automorphism group of the Lie bialgebra and that of the associated graded are given by the KV and KRV groups. This section summarises definitions of these spaces and results in [AKKN23].
@@
  • $\hat{T}(H)$ is regarded as an $\hat{L}(H)$-module by the left multiplication. For $w=x_{i},y_{i},z_{j}$, $d_{w}\colon\hat{L}(H)\to\hat{T}(H)$ is a continuous Lie algebra 11 11-cocycle specified by $d_{w}(w^{\prime})=\delta_{ww^{\prime}}$ for $w^{\prime}=x_{i},y_{i},z_{j}$ using Kronecker’s delta.
@@
    $\operatorname{div}_{x,y,z}\colon\operatorname{tDer}^{+}(\hat{L}(H))\to\operatorname{Der}(\hat{L}(H))\xrightarrow{\operatorname{div}_{x,y,z}}|\hat{T}(H)|.$
@@ 4.1 Tangential Derivations
For the definition of a derivation on a $\mathbb{K}$-linear category, see Section 4 of [Tan24a], for example.
@@
Proof. Suppose we are given a tangential derivation $(u,u_{1},\dotsc,u_{n})$. We construct a $\mathbb{K}$-linear category derivation $f$. For a loop $\alpha$ based at $*_{0}$, we set $f(\alpha)=u(\alpha)$. For a path $\delta$ from $*_{j}$ to $*_{0}$ such that $\delta\gamma_{j}\delta^{-1}$ is the $j$-th boundary loop $\partial_{j}\Sigma$, we set $\delta^{-1}f(\delta)=u_{j}$. This is well-defined: for another such path $\delta^{\prime}$, the loop $\delta^{\prime-1}\delta$ is based at $*_{0}$ and we have $(\delta^{\prime-1}\delta)\gamma_{j}=\gamma_{j}(\delta^{\prime-1}\delta)$. Since $\mathscr{G}$ is a free groupoid, we have $\delta^{\prime-1}\delta=\gamma_{j}^{-m}$ for some $m\in\mathbb{Z}$. We compute
@@ 4.2 Connections on Modules
The $\mathbb{K}$-Hopf groupoid structure of the groupoid algebra $\mathbb{K}\mathscr{G}$ is defined by the following: a counital coassociative (cocommutative) coalgebra structure on the Hom-spaces given by, for $v,w\in\mathrm{Ob}(\mathscr{G})=V$ and $x\in\mathscr{G}(v,w)$,
@@
The completion $\widehat{\mathbb{K}\mathscr{G}}$ with respect to the filtration in Definition 4.7 and its associated graded $\hat{\mathscr{T}}$ are automatically (complete) Hopf groupoids. We set
@@
Recall that the module over an algebra is said to be dualisable if it is finitely generated and projective.
@@
This decomposition is given by the free groupoid version of the Fox derivatives. Then, since the $\mathbb{K}\mathscr{G}$-modules $\mathbb{K}\mathscr{G}(\,\cdot\,,*_{0})$ and $\mathbb{K}\mathscr{G}(\,\cdot\,,*_{j})$ are projective and finitely generated (in fact, they are generated by a singe element), the claim follows. For analogous statements and their proofs, see Proposition 3.8 in [Tan24b] and Lemma 5.6 in [Tan24a].∎
@@
Take the free-generating system $\mathcal{C}$ of $\mathscr{G}$ as in Figure 2 of [Tan24a]. We define the connection $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}\colon\mathscr{N}\to\Omega^{1}\mathbb{K}\mathscr{G}\otimes_{\mathbb{K}\mathscr{G}}\mathscr{N}$ by
@@
Here, $a_{i}$, $b_{i}$ and $c_{j}$ are the rotation numbers of the generators $\alpha_{i},\beta_{i},\gamma_{j}\in\mathcal{C}$. Extend $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ to the completion $\hat{\mathscr{N}}$ by continuity, which is a $\widehat{\mathbb{K}\mathscr{G}}$-module. A derivation action of $f\in\operatorname{Der}_{\mathrm{Hopf}}(\widehat{\mathbb{K}\mathscr{G}})$ on the module $\hat{\mathscr{N}}$ is given by the restriction $f|_{\operatorname{Ker}(s)}$.
@@
Let $\hat{\mathscr{T}}$ as in Definition 4.7 and $\hat{\mathscr{N}}_{\mathrm{gr}}=\operatorname{Ker}(s\colon\hat{\mathscr{T}}\to\mathbb{K}V)$, which is also dualisable over $\hat{\mathscr{T}}$ since it is isomorphic to $\mathscr{N}$ via $\theta_{\mathrm{exp}}$. We define the connection $\nabla^{\prime}\!\!_{H,\mathsf{fr}}$ on $\hat{\mathscr{N}}_{\mathrm{gr}}$ by
@@
This is an extension of the flat connection $\nabla^{\prime}\!\!_{z}$ defined in Section 6 of [Tan24b]. Also this is the associated graded of $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ above since $(d\beta_{i})\beta_{i}^{-1}$ and $(d\alpha_{i})\alpha_{i}^{-1}$ in the definition of $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ have the weight $\geq 1$.
@@
Before the proof, we look at the counterpart of the connection $\nabla^{\prime}\!_{\mathcal{C},\mathsf{fr}}$ on a $\mathbb{K}\mathscr{G}$-bimodule $\Omega^{1}\mathbb{K}\mathscr{G}$. For a $\mathbb{K}$-algebra $A$, we denote the copy of an element $a\in A$ in $A^{\mathrm{op}}$ by $\bar{a}$. We denote by $A^{\mathrm{e}}$ the enveloping algebra of $A$.
@@
We define the connection $\nabla\!_{\mathcal{C},\mathsf{fr}}\colon\Omega^{1}\mathbb{K}\mathscr{G}\to\Omega^{1}\mathbb{K}\mathscr{G}^{\mathrm{e}}\otimes_{\mathbb{K}\mathscr{G}^{\mathrm{e}}}\Omega^{1}\mathbb{K}\mathscr{G}$ on the $\mathbb{K}\mathscr{G}$-bimodule $\Omega^{1}\mathbb{K}\mathscr{G}$ by
@@
First, the source map corresponds to the left $A$-module map $s\colon A\to\mathbb{K}V$, where the corresponding $A$-module to $\mathbb{K}V$ is denoted again by $\mathbb{K}V$. In addition, $\mathscr{N}$ corresponds to $N=\operatorname{Ker}(s\colon A\to\mathbb{K}V)$. Next, consider the functor